OpenStudy (hanner_b_nanner):

Does anybody wanna help me with some Algebra 2??? Fan&&Medal!! :D

OpenStudy (anonymous):

me !!!

OpenStudy (anonymous):

Sure why not

OpenStudy (anonymous):

problem?

OpenStudy (hanner_b_nanner):

okeee dokeee!!! Use the four functions below for this question: f(x) is attached below... g(x): g(x) = 3x − 4 h(x): g(x) = 3x − 4 j(x): x j(x) ______ -2 2 2 10 6 22 Compare and contrast these four functions using complete sentences. Focus on the slope and y-intercept first, and then on any additional properties of each function.

OpenStudy (hanner_b_nanner):

grrr... h(x) is Kelly had 4 apples in her kitchen on Monday. On Tuesday, only 1 was left. ** sorry!

OpenStudy (hanner_b_nanner):

andddd g(x) = 3x - 4**

OpenStudy (anonymous):

so the equations are g(x)=3x -4 h(x)= -3x + 7 and j(x)=3x + 4 all in order. they all have the same slope except h(x) is flipped. because it's negative it's going from top left down to right bottom. the other two are positive slopes so its from bottom left to top right. j(x) has a y intercept of 4 and it starts in the middle between g(x) and h(x). h(x) has the highest y intercept and g(x) has the lowest

OpenStudy (hanner_b_nanner):

oh my goodness, thank you SO much!!!!!!!!! :D

OpenStudy (anonymous):

your welcome